Mark wants to create several passwords for several accounts. Each password is 9 characters consisting of arangements of the 4 letters of his name followed by 5 digits from his birthdate 8/21/73. How many passwords can he create ?
Answer by jim_thompson5910(35256)

There are 4! = 4*3*2*1 = 24 ways to arrange the letters of the name "mark"

There are 5! = 5*4*3*2*1 = 120 ways to arrange the digits: 8, 2, 1, 7, 3

So there are 24*120 = 2,880 different passwords

This may sound like a lot, but to a computer this is nothing. A computer can try thousands of passwords a second. So Mark is better off using all the letters of the alphabet (both lower and upper case) along with all ten numeral digits to get a really safe password.